OpenStudy (linda):

solve for x: log 4 (x^2+3x) - log 4 (x+5) = 1 plzz hellp

OpenStudy (anonymous):

Log(a)-Log(b) = Log(a/b) log4 (x^2+3x) - log4 (x+5) =1 log4 (x^2+3x)/ (x+5) = 1 (x^2+3x)/(x+5) = 4^1 x^2+3x=4(x+5) x^2+3x=4x+20 x^2+3x-4x-20=0 x^2-x-20=0 x^2-5x+4x-20=0 x(x-5)+4(x-5)=0 (x-5)(x+4)=0 x=5 or x=-4

OpenStudy (linda):

thanks!!

OpenStudy (anonymous):

You are welcome